# Graph Memory Learning: Imitating Lifelong Remembering and Forgetting of Brain Networks

### v1

@@ IV-C MGHPL: multi-granular hierarchical progressive learning mechanism
Brain-Inspiration. The brain’s cognition process of knowledge involves the coordination and hierarchical progressive processing mechanisms of the partitions of the cerebral cortex. When a piece of knowledge enters, the corresponding primary sensory area of the cerebral cortex responds. Due to limitations, not all extracted features from the area can provide the fine information required for an accurate judgment. Therefore, for primary areas of the cerebral cortex that perceive blur, the brain will pass the information to the corresponding secondary partitions. In the secondary partition, insufficient cognition in primary perceptual areas will be targeted and supplemented. When participating in decision-making, the information extracted by each cerebral cortex is fused to form higher-level cognitive features. From primary perception to secondary processing to higher-level cognition, the brain’s process of learning knowledge is progressive and in-depth.
Here, we still use the example of distinguishing tangerines and oranges. When information such as the color and taste of fruits is transmitted to the brain, it is sent to different cortical areas by the sensory system. For example, visual signals are sent to the occipital lobe, taste signals are sent to the temporal lobe, etc. The flavors of the two fruits are relatively easy to distinguish, with tangerines being sweeter and orange being sweet with a hint of sourness. Therefore, primary perception and judgment in the temporal lobe area are relatively accurate and do not require the assistance of secondary partitions. Visually, however, they are both orange and difficult to distinguish. Information about the fruit’s skin characteristics will need to be sent to the subdivision to aid in judgment. Ultimately, the information processed by these different cortical areas undergoes complex processing and fusion to help us make the correct cognitions and decisions.
First FGGL. Inspired by the above, this paper proposes the MGHPL mechanism based on feature graph grain learning (FGGL), as shown in Fig. 2. This mechanism takes the first-level graph shards and the second-level graph shards of the static graph at time $t_{0}$ given by the MGP algorithm as input. The first is the first-level FGGL. It is important that FGGL be related to subsequent continuous streaming graph learning as well as the response to node data IR and FR. This part of the idea refers to Wang et al.’s [14] implicit encoding of graph structure into the node feature graph, which breaks the shackles of node learning being restricted by the graph structure. This idea of converting a single node into a feature graph and converting the graph into a collection of node feature graphs (i.e., graph grains) inspires us. Specifically, the features of $F$ nodes $\left\{f_{1},f_{2},\dots,f_{F}\right\}$ in the feature graph of node $v_{i}$ are obtained by splitting the original feature $\mathbf{X}_{i}\in\mathbb{R}^{F\times C}$. The construction of the adjacency matrix in the feature graph can be expressed by the following formula:
@@
where $\mathrm{sgnroot}(x)=\mathrm{sign}(x)\sqrt{\left|x\right|}$. For $\forall c\in\left\{1,2,\dots,C\right\}$, $\mathbf{X}_{i}^{c}\in\mathbb{R}^{F}$ represents all features in the $c$-th channel of node $v_{i}$. $\omega_{i,j}^{c}$ represents the weight of the $c$-th channel edge $e_{i,j}$. $\mathbf{w}_{i},\mathbf{w}_{j}\in\mathbb{R}^{F}$, $q\in\mathbb{R}$ are learnable attention parameters. At this point, all first-level graph shards are converted into graph grains according to Eq.4. To avoid confusion, graph grains and graph shards share the same symbol:
@@ IV-D Memory Forgetting Module
Brain-Inspiration. Human cognition of knowledge is a process of continuous adjustment. During continuous learning, some information may negatively impact cognitive activities, endangering the overall cognitive system. The brain must correct errors promptly. Additionally, for information that the brain subjectively deems unnecessary to retain, it initiates a forgetting process as needed. Biologically, this forgetting operation involves mechanisms like autonomous synaptic remodeling of neurons in relevant partitions.
Let us consider guitar practice as an example. When encountering incorrect fingering, or uninteresting skills, the brain initiates the forgetting process to correct these errors. This occurs because the brain has a mechanism to eliminate invalid or erroneous information. During forgetting, the relevant neural connections and synapses are readjusted. Concurrently, the brain releases more space and resources, allowing for more efficient learning and mastery of the correct skills.
@@
$\left\{\begin{array}[]{l}\mathcal{S}_{i}\leftarrow\mathcal{S}_{i}\setminus\left\{v_{\mathrm{forget}},e_{\mathrm{forget},u}\mid\forall v_{u}\in\mathcal{N}(v_{\mathrm{forget}})\right\}\\ \mathcal{P}_{j}^{i}\leftarrow\mathcal{P}_{j}^{i}\setminus\left\{v_{\mathrm{forget}},e_{\mathrm{forget},u}\mid\forall v_{u}\in\mathcal{N}(v_{\mathrm{forget}})\right\},\end{array}\right.$ (10)
@@
Re-FGGL. Based on the updated graph shards of Eq.10 and Eq.4, the graph grains have been updated in preparation for further learning. The irrelevant graph grains remain unaffected. For specific forgetting measures, this paper employs the scratch method. This approach ensures the complete erasure of all lingering memories of forgotten data within the system. Importantly, this scratch process is resource-efficient, as it solely involves pertinent sub-models, leaving the rest unaffected. Once the relevant sub-models have undergone retraining, they can seamlessly integrate into the decision-making process, serving as fresh references, thereby enhancing convenience.
@@ IV-E Memory Remembering Module
Brain-Inspiration. As humans grow, Brain knowledge is constantly being updated. Biologically, the brain’s memory adjustment function is autonomous. New information from sensory organs is adaptively transmitted to the most relevant cortical areas to form a response. In these specific brain regions, neurons encode and decode new sensory inputs, comparing and integrating them with existing memories and knowledge. To avoid conflicts between new information and existing knowledge, the brain may selectively build neural connections to introduce new information into memory.
Consider a beginner learning guitar. Initially, the beginner knows only the style and playing method of the guitar. During practice, the sensory system captures information such as chords, techniques, and music scores. When encountering a new chord, the auditory sense converts the stimulus into a signal transmitted by neurons in the temporal cortex. Eventually, the brain forms a memory and associates it with previous chord experiences.
@@ V-A Datasets and Baselines
Datasets. Here we evaluate the advantages, functionality, effectiveness, and rationale of the BGML framework across five commonly utilized node classification datasets [52, 53]. The datasets comprise Cora, CiteSeer, PubMed, Coauthor-CS, and Coauthor-Physics. Detailed information regarding these datasets is provided in Table I. The rest of the information about these datasets is presented in Appendix B.
@@
Dataset Cora CiteSeer PubMed Coauthor-CS Coauthor-Physics
Nodes 2708 3327 19717 18333 34493
Edges 5429 4732 44338 81894 247962
Features 1433 3703 500 6805 8415
Classes 7 6 3 15 5
Baselines. Considering the novelty and complexity of graph memory learning, several factors were taken into account when choosing the baseline method for comparative experiments. For comparison, we chose two types of methods: the first includes classic and widely-used methods in the current graph field, such as MLP, ChebNet [54], GCN [23], GraphSAGE [25], GAT [24], and APPNP [55]; the second comprises methods related to graph memory learning, namely the graph lifelong learning leader method FGNs [14] and the graph unlearning pioneer GraphEraser [16].
The choice was made as GML is a novel subject that is only first discussed in this work, and no previous works have been found to serve as baselines for comparative experiments. The GML described in Definition 1 must, in essence, be capable of machine unlearning, continuous learning, and strong graph representation at all times. As a result, we chose the classic work in the aforementioned graph learning field that is most pertinent to these capability requirements for conducting comparative experiments to test model capabilities. The most recent works of GLL and GUL were not chosen as baselines for two reasons. On the one hand, FGNs and GraphEraser are still very competitive in current research. On the other hand, many GLL and GUL methods are not well-suited to GML problems due to their disparate designs.
@@ V-E Evaluation of Incremental Learning Efficacy
V-F Ablation Studies
@@
TABLE VII: Ablation studies. $\surd$” indicates that the component is included in the model, whereas “$\times$” indicates that it is not included.
@@
V-G Analysis of FR and IR Quantities
@@ VI Conclusion
In response to the challenge of continuous data influx and withdrawal in graph machine learning, this paper highlights the pressing issues of graph memory learning for the first time. Thus, inspired by the modular and hierarchical structure of the brain, as well as the brain network dynamics, the paper presents a general GML framework. BGML adeptly addresses diverse memory challenges in dynamic evolution scenarios and embodies the principle of selective remembering and forgetting. Experimentally, BGML outperforms other baseline models to varying degrees across different learning and memory tasks on the five datasets.
@@
However, we also identified certain issues that warrant further exploration in the future. Specifically, (1) there exist interareal communication subspaces, which enhance learning. How can we establish connections between BGML’s sub-models without compromising its efficient forgetting ability? (2) BGML sequentially processes FR and IR within the same time stamp. Is it possible to cleverly overwrite or modify the information slated for forgetting using IR, thereby offering a more uniform approach to handling memory issues?
Appendix A The Algorithm Framework of BGML
Algorithm 1 The algorithm for BGML.
1:  for $\mathrm{timestamp}=t_{0}$ to $t_{t}$ do
2:     if $\mathrm{timestamp}=t_{0}$ then
3:        Perform MGP (Eq.1) on $\mathcal{G}_{0}$ (Eq.2 or Eq.3).
4:        First FGGL (Eq.4, Eq.5, and Eq.6).
5:        Perform progress-aware module (Eq.7).
6:        Second FGGL (Eq.8).
8:     else
9:        Receive FR $\mathcal{V}_{\mathrm{delete}}$, update MGP(Eq.10).
10:        Perform ForgetGuide Module (Eq.11).
11:        Re-FGGL(Eq.4, Eq.6, Eq.8).
13:        Incremental FGGL (Eq.6, Eq.8).
15:     end if
16:     Update timestamp $t_{i}\to t_{i+1}$.
17:  end for
Firstly, at time $t_{0}$, the graph $\mathcal{G}_{0}$ is segmented into multiple shards using the balanced graph partitioning methods BLPA or BEKM. Secondly, multi-granular graph shards are fed into MGHPL based on FGGL for learning. Specifically, this paper transforms nodes into feature graphs by constructing a feature graph adjacency matrix, thereby converting graph shards into multi-granular graph grains. Subsequently, first-level graph grains are employed to train the first-level sub-model constructed by the feature graph networks. Moreover, these sub-models undergo evaluation and scoring. Based on the scores, the first-level sub-models requiring support can be progressively perceived, and corresponding second-level graph grain learning can be executed. Finally, the result $\chi_{\mathrm{final}}^{0}$ of task $\mathrm{T}_{0}$ can be derived by integrating sub-models following the hierarchical progressive strategy.
Appendix B Details of the datasets
Appendix C Datasets Division for Tasks and Implementation
Datasets Division for Tasks. Initially, we utilized a random division method to split all datasets into separate training and test sets in an 8:2 ratio. For regular graph learning and graph unlearning tasks, all training graphs were directly fed into the model for training. However, in scenarios involving data-incremental and class-incremental, as well as graph memory issues, the training set was divided into an initial training graph (at time $t_{0}$) and random incremental data input through streaming. The division ratio and the number of incremental data per time step can be adjusted flexibly based on task requirements (Appeddix D). Similarly, the number of FR can be adjusted accordingly (Appeddix D).
Appendix D Hyperparameters Analysis
(a) Number of first-level graph shards
(b) Number of second-level graph shards
Figure 5: The impact of $k$ and $l$.
(a) Number of sub-models retrieved
(b) Number of self-selected neighbors
Figure 6: The impact of $\tau$ and $\mu$.
[Uncaptioned image] Jiaxing Miao received his M.S. degree from the School of Sciences, China Jiliang University, Hangzhou, China, in 2023. He is currently pursuing a Ph.D. degree from the College of Electronics and Information Engineering, Tongji University, Shanghai, China. His current research interests mainly include graph representation learning, lifelong learning, and machine unlearning.
[Uncaptioned image] Liang Hu (@M in 2022) is a full professor with Tongji University. His research interests include recommender systems, machine learning, data science and interdisciplinary intelligence. He has published a number of papers in top-rank international conferences and journals, including WWW, IJCAI, AAAI, ICDM, NeurIPS, TOIS, TNNLS, TKDE. He has been invited as the program committee of more than 50 top-rank AI international conferences, including AAAI, IJCAI, ICDM, CIKM, CVPR, and KDD.
[Uncaptioned image] Qi Zhang received the PhD degree from the Beijing Institute of Technology, Beijing, China, and the University of Technology Sydney, Sydney, NSW, Australia, in 2020. He is currently a Research Fellow at Tongji University, Shanghai, China. He has authored high-quality papers in premier conferences and journals, including AAAI, IJCAI, SIGIR, TKDE, TNNLS, and TOIS. His research interests include collaborative filtering, sequential recommendation, learning to hash, and MTS analysis.
[Uncaptioned image] Longbing Cao (@SM in 2006) received the PhD degree in pattern recognition and intelligent systems from the Chinese Academy of Science, China, and the PhD degree in computing sciences from the University of Technology Sydney, Australia. He is a distinguished chair professor with Macquarie University, an ARC future fellow (professorial level), and the EiCs of IEEE Intelligent Systems and J. Data Science and Analytics. His research interests include AI, data science, machine learning, behavior informatics, and their enterprise applications.

### v2

@@ IV-C MGHPL: multi-granular hierarchical progressive learning mechanism
To address the frequent changes in local graph structures and mitigate the conflict between memory and forgetting, this paper, inspired by brain network principles such as the coordination and hierarchical processing of cortical regions, brain information fusion and decision-making (Remark 2), proposes a multi-granular hierarchical progressive learning mechanism based on the decoupling of graph knowledge.
Remark 2.
The Coordination and Hierarchical Processing of Cortical Regions refers to the coordinated functioning of different cortical areas during cognitive processes, progressing from primary sensory regions to secondary processing areas, and ultimately to higher-level cognitive integration and decision-making. This reflects the gradual deepening of the brain’s cognitive processes. The Brain Information Fusion and Decision-Making refers to the extraction of information by modular and multi-level cortical regions, which is then fused and integrated to form higher-level cognitive features. It is crucial for accurate judgment and decision-making.
First FGGL. Inspired by the above, this paper proposes the MGHPL mechanism based on feature graph grain learning (FGGL), as shown in Fig. 2. This mechanism takes the first-level graph shards and the second-level graph shards of the static graph at time $t_{0}$ given by the MGP algorithm as input. The first is the first-level FGGL. It is important that FGGL be related to subsequent continuous streaming graph learning as well as the response to node data IR and FR. This part of the idea refers to Wang et al.’s [13] implicit encoding of graph structure into the node feature graph, which breaks the shackles of node learning being restricted by the graph structure. Based on this, we also convert a single node into a feature graph and converting the graph into a collection of node feature graphs (i.e., graph grains). The adjacency matrix of the feature graph is constructed as follows:
@@
where $\mathrm{sgnroot}(x)=\mathrm{sign}(x)\sqrt{\left|x\right|}$ and $\mathbf{X}_{i}\in\mathbb{R}^{F\times C}$. For $\forall c\in\left\{1,2,\dots,C\right\}$, $\mathbf{X}_{i}^{c}\in\mathbb{R}^{F}$ represents all features in the $c$-th channel of node $v_{i}$. $\omega_{i,j}^{c}$ represents the weight of the $c$-th channel edge $e_{i,j}$. $\mathbf{w}_{i},\mathbf{w}_{j}\in\mathbb{R}^{F}$, $q\in\mathbb{R}$ are learnable attention parameters. At this point, all first-level graph shards are converted into graph grains according to Eq.4. To avoid confusion, graph grains and graph shards share the same symbol:
@@
The innovation of MGHPL lies in the gradual adaptation of the learning granularity, which mirrors the human cognitive process. Progress-aware module is the key technology for achieving this goal, which evaluates and provides real-time feedback on the cognitive ability of each sub-model, dynamically adjusting the integration framework to facilitate the incremental supplementation and integration of cognitive information at different granularities. This progressive, dynamically adjustable multi-granularity learning system enhances the model’s ability to accurately delineate the boundary between retained and forgotten knowledge.
@@ IV-D Memory Forgetting Module
The principles of autonomous synaptic selection, synaptic remodeling, and neural resource allocation in the brain’s information forgetting (Remark 3) have inspired our approach to handling continuous forgetting request inputs. The memory forgetting module, driven by the forgetting-guided module, implements an efficient and localized pinpoint clearing model.
Remark 3.
Autonomous Synaptic Selection refers to the brain’s ability to automatically identify and adjust the corresponding synaptic connections when confronted with erroneous information. Synaptic Remodeling involves the reorganization or restructuring of neural synapses in relevant regions during the forgetting process. This adjustment enables accurate forgetting. Neural Resource Allocation refers to the redistribution of neural resources during forgetting or synaptic remodeling, allowing the brain to focus on processing more meaningful information or skill acquisition, thereby enhancing learning efficiency.
@@
$\left\{\begin{array}[]{l}\mathcal{S}_{i}\leftarrow\mathcal{S}_{i}\setminus\left\{v_{\mathrm{forget}},e_{\mathrm{forget}}\right\}\\ \mathcal{P}_{j}^{i}\leftarrow\mathcal{P}_{j}^{i}\setminus\left\{v_{\mathrm{forget}},e_{\mathrm{forget}}\right\},\end{array}\right$. (10)
@@
Re-FGGL. Based on the updated graph shards of Eq.10 and Eq.4, the graph grains have been updated in preparation for further learning. The irrelevant graph grains remain unaffected. For specific forgetting measures, this paper employs the scratch method. It solely involves pertinent sub-models, leaving the rest unaffected. Once the relevant sub-models have undergone retraining, they can seamlessly integrate into the decision-making process, serving as fresh references, thereby enhancing convenience.
@@ IV-E Memory Remembering Module
The process of integrating new information with existing long-term memory involves not only the storage of information but also its reinforcement and consolidation. To address the conflicts arising from the low credibility of new knowledge structures, this paper, inspired by principles in brain dynamics such as cross-temporal memory comparison and integration, selective neural connection construction, and experience association (Remark 4), proposes the construction of a memory remembering module centered around ISAO. This module organizes reliable graph structure and buffers the conflicts caused by memory and forgetting.
Remark 4.
Cross-temporal Memory Comparison and Integration refers to the process of comparing current information with historical memories, integrating it based on the assessment of its variability or consistency. This process involves functional coordination between the prefrontal cortex and hippocampus. Selective Neural Connection Formation refers to the brain’s ability to selectively establish new neural connections to avoid conflicts between new information and existing knowledge, facilitating the incorporation of new information into the memory system. Experience Association is the process through which the brain links and integrates new experiences with past ones, forming memories. This process aids in transforming new skills and information into long-term memory, enhancing their application and recognition.
@@
ISAO simulates brain network dynamics to adaptively attribute and update knowledge, improving information reliability and consistency. Its key innovation is that new nodes select graph grains based on their features and adaptively adjust neighbor relationships, using the most relevant historical experiences to promote rapid information fusion and updating. This mechanism mimics how the human brain integrates new and old knowledge. The approach ensures the credibility of local graphs and neighboring nodes, while also adaptively aligning memory and unlearning requests with the corresponding sub-models of MGHPL, effectively mitigating the conflict between memory and forgetting.
@@
IV-F The Algorithm Framework of BGML
Algorithm 1 The algorithm for BGML.
1:for $\mathrm{timestamp}=t_{0}$ to $t_{t}$ do
2:  if $\mathrm{timestamp}=t_{0}$ then
3:   Perform MGP (Eq.1) on $\mathcal{G}_{0}$ (Eq.2 or Eq.3).
4:   First FGGL (Eq.4, Eq.5, and Eq.6).
5:   Perform progress-aware module (Eq.7).
6:   Second FGGL (Eq.8).
8:  else
9:   Receive FR $\mathcal{V}_{\mathrm{delete}}$, update MGP(Eq.10).
10:   Perform ForgetGuide Module (Eq.11).
11:   Re-FGGL(Eq.4, Eq.6, Eq.8).
13:   Incremental FGGL (Eq.6, Eq.8).
15:  end if
16:  Update timestamp $t_{i}\to t_{i+1}$.
17:end for
Firstly, at time $t_{0}$, the graph $\mathcal{G}_{0}$ is segmented into multiple shards using the balanced graph partitioning methods BLPA or BEKM. Secondly, multi-granular graph shards are fed into MGHPL based on progress-aware module for learning. Specifically, this paper transforms nodes into feature graphs by constructing a feature graph adjacency matrix, thereby converting graph shards into multi-granular graph grains. Subsequently, first-level graph grains are employed to train the first-level sub-model constructed by the feature graph networks. Moreover, these sub-models undergo evaluation and scoring. Based on the scores, the first-level sub-models requiring support can be progressively perceived, and corresponding second-level graph grain learning can be executed. Finally, the result $\chi_{\mathrm{final}}^{0}$ of task $\mathrm{T}_{0}$ can be derived by integrating sub-models following the hierarchical progressive strategy.
@@ V-A Datasets and Baselines
Datasets. Here we evaluate the advantages, functionality, effectiveness, and rationale of the BGML framework across nine node classification datasets. These datasets include widely-used benchmark datasets—Cora, CiteSeer, PubMed, Coauthor-CS, and Coauthor-Physics [59, 60]—as well as real-world datasets, namely Ogbn-arXiv [13], Reddit [29], Flickr [61] and Yelp [61]. Detailed information regarding these datasets is provided in Table I. The rest of the information about these datasets is presented as follows:
@@
Dataset Nodes Edges Features Classes
Cora 2708 5429 1433 7
CiteSeer 3327 4732 3703 6
PubMed 19717 44338 500 3
Coauthor-CS 18333 81894 6805 15
Coauthor-Physics 34493 247962 8415 5
Ogbn-arXiv 169343 1166243 128 40
Reddit 232965 11606919 602 41
Flickr 89250 899,756 500 7
Yelp 716847 6977410 300 100
  • Real-world datasets: Ogbn-arXiv [13], Reddit [29], Flickr [61] and Yelp [61]. Both datasets are sourced from real-world scenarios and serve as representative examples of large graph datasets. The Ogbn-arXiv dataset represents the citation network of all computer science-related papers on the arXiv platform. Notably, the Reddit dataset is a dynamic graph derived from the social news website Reddit, encompassing users from various communities and their interactions. In the Flickr network, nodes represent individual images. An edge is created between two images if they share common characteristics, such as the same geographic location or gallery. The Yelp dataset contains a recommendation network, where an edge signifies that the connected users are friends.
Baselines. Considering the novelty and complexity of graph memory learning, several factors were taken into account when choosing the baseline method for comparative experiments. For comparison, we selected two categories of methods. The first includes classic and widely-used methods in the current graph field, such as MLP, ChebNet [62], GCN [27], GraphSAGE [29], GAT [28], APPNP [63], and Mixup [61]. It is worth noting that when GCN and GAT are inconvenient to use with large-scale datasets, we opt to combine them with GraphSAINT [64] before making comparisons. The second category comprises methods related to graph memory learning, namely the graph lifelong learning leader methods: FGNs [13], SEA-ER [16] SEM [50], OTGNet [49], TACO [44], and the graph unlearning pioneer methods: GraphEraser [17], SGC-Unlearn [18], GraphDelete [57].
The choice was made as GML is a novel subject that is only first discussed in this work, and no previous works have been found to serve as baselines for comparative experiments. The GML described in Definition 1 must, in essence, be capable of machine unlearning, continuous learning, and strong graph representation at all times. As a result, we chose the classic work in the aforementioned graph learning field that is most pertinent to these capability requirements for conducting comparative experiments to test model capabilities. We selected competitive methods from the domains of GUL and GLL for comparative experiments, based on the specific adaptability requirements of different tasks. Unfortunately, methods that are not suitable for GML problems could not be included in the comparison. Notely, the model-agnostic graph methods that appear in this paper, such as GraphEraser, are all based on GCN to form a specific model.
@@ V-B Experiment Settings
Datasets Division for Tasks. Initially, we utilized a random division method to split all datasets into separate training and test sets in an 0.8:0.2 ratio. For regular graph learning and graph unlearning tasks, all training graphs were directly fed into the model for training. However, in scenarios involving data-incremental and class-incremental, as well as graph memory issues, the training set was divided into an initial training graph (at time $t_{0}$) and random incremental data input through streaming. The division ratio and the number of incremental data per time step can be adjusted flexibly based on task requirements (V-I). Similarly, the number of FR can be adjusted accordingly (V-I). Specifically, Fig. 4 presents the dynamic experiments on the applicability of BGML and various baselines to the GML problem across real-world datasets. The experiments include the large-scale graph dataset Ogbn-Arxiv, as well as three dynamic networks with natural temporal information (Reddit, Flickr, and Yelp). The partitioning of Ogbn-Arxiv follows the same setup as other commonly used node classification datasets. For the remaining three dynamic networks with timelines, we select the midpoint of the overall timeline as the $t_{0}$ moment for the GML problem. The subsequent incremental nodes and the dynamic changes in the graph correspond to the natural temporal evolution of the dataset, with forgotten nodes being randomly selected. The number of FR can be adjusted flexibly, for example, 50 or 100 nodes per round. For clarity, the line graph in Fig. 4 reports the performance at $t_{0}$ and at four equidistant time nodes ($t_{1}$ to $t_{4}$) on the remaining timeline.
@@ V-E Evaluation of Incremental Learning Efficacy
V-F Evaluation on Real-world Datasets
In Fig.4, we present a dynamic visualization of the performance trends of BGML and several baseline models in solving GML problems across four real-world datasets using a timeline and line graph. The main goal of this experiment is to assess the advantages and unique capabilities of BGML in addressing the challenges of GML within large-scale, real-world dynamic networks that inherently include temporal information.
The experimental results provide compelling evidence that, across both datasets, BGML outperforms other models in terms of accuracy and stability throughout most periods of the GML task. This highlights BGML’s strong capacity to handle GML problems continuously. More notably, when compared to traditional graph learning models and other baseline methods, BGML’s line graph shows a more stable trend, thus demonstrating its robustness in dynamic, evolving environments. The findings further support the central argument of this paper: BGML effectively mitigates the conflict between ”memory” and ”forgetting.” In contrast, other benchmark models suffer from significant performance fluctuations and degradation due to the frequent introduction and removal of new and old data. By incorporating MGHPL and ISAO, BGML efficiently processes incremental updates, maintains robust memory retention of historical information, while selectively forgetting certain details, even in the context of a constantly evolving graph.
@@ V-F Evaluation on Real-world Datasets
(a) Arxiv
(b) Reddit
(c) Flickr
(d) Yelp
Figure 4: Performance of different models at different time steps in GML.
V-G Ablation Studies
@@
TABLE VII: Ablation studies. $\surd$” indicates that the component is included in the model, whereas $\times$” indicates that it is not included.
@@
V-H Hyperparameters Analysis
(a) Number of first-level graph shards
(b) Number of second-level graph shards
Figure 5: The impact of $k$ and $l$.
(a) Number of sub-models retrieved
(b) Number of self-selected neighbors
Figure 6: The impact of $\tau$ and $\mu$.
V-I Analysis of FR and IR Quantities
@@ VI Conclusion and Discussion
In response to the challenge of continuous data influx and withdrawal in graph machine learning, this paper highlights the pressing issues of graph memory learning for the first time. Thus, inspired by the modular and hierarchical structure of the brain, as well as the brain network dynamics, the paper presents a general GML framework. BGML adeptly addresses diverse memory challenges in dynamic evolution scenarios and embodies the principle of selective remembering and forgetting. Experimentally, BGML outperforms other baseline models to varying degrees across different learning and memory tasks on the nine datasets.
@@
Limitation and Future Work. This study uses nodes as instances study to investigate strategies for handling the influx and withdrawal of nodes in dynamic graphs. In practical scenarios, edge evolution is also prevalent. In the future, Our research will address edge-level tasks in GML problems. Additionally, during the course of this research, potential directions for future exploration emerged, particularly in mitigating the conflict between memory and forgetting. Specifically, (1) whether the experiences of individual sub-models in processing evolution requests for similar structures can be stored and leveraged. there exist interareal communication subspaces in the brain, which enhance learning. How can we establish connections between BGML’s sub-models without compromising its efficient ability? (2) The conflicts stem from the need to distinguish between memory and forgetting requests. BGML sequentially processes FR and IR within the same time stamp. Is it possible to cleverly overwrite or modify the information slated for forgetting using IR, thereby offering a more uniform approach to handling memory issues?